OpenStudy (anonymous):

(2y)/(y+3)+(6)/(y+3)

OpenStudy (anonymous):

2y/y+3 + 6/y+3 =2y+6 / 2(y+3) GCF of the numerator =2(y+3) / 2(y+3) =1